An old wheat-grinding wheel in a museum actually works. The sign on the wall says that the wheel has a rotational acceleration of 210 rad/s^2 as it's spinning rotational speed increases from zero to 1900 rpm. How long does it take the wheel to attain this rotational speed?
Answer #1

Rotational acceleration of the wheel is, \alpha = 210 rad/s2

Initial spinning rotational speed of the wheel is, \omega _{i} = 0 rpm

1 rpm = (1/60) revolution per second = 2\pi * (1/60) rad/s

Thus,

\omega _{i} = 0 rad/s

Final spinning rotational speed of the wheel is, \omega _{f} = 1900 rpm = 198.99 rad/s

Now, from the equations of motion (for rotational motions),

\omega _{f} = \omega _{i} + \alpha t

where t is the time taken by the wheel to attain its final rotational speed.

198.99 = 0 + 210 * t

t = 198.99 / 210

t = 0.947571429 s = 0.95 s